# Training Neural Networks on Data Sources with Unknown Reliability

### v2

@@ 3 Methods
Figure 1: Visualisation of Equation 5. Each colour represents the loss values from a single source over a small number of steps, with its density weighted by its temperature, $w(s)={1}/{f(C_{s})}$. This shows how sources contribute to $\mu_{s^{\prime}}$ and $\sigma_{s^{\prime}}^{2}$ as their $C_{s}$ changes during training and given the leniency $\lambda$. These values are synthetic and for demonstration.
Intuitions
@@
Figure 1 shows how the training process evolves over time when using LAP. At first, all sources are considered equally reliable and so the weighted mean of the loss values is the mean of all sources. As the source temperature changes, the weighed mean plus $\lambda$ standard deviations of the losses moves towards the sources with lower expected losses, allowing for more learning from the these over time.
@@
Figure 2 illustrates how $1-d_{s}$ (reciprocal of the temperature) in Equation 6 develops during training for each source and different $H$, $\delta$, and $\lambda$ values. The first row shows how varying the depression strength, $\delta$ affects the rate at which the learning on a source is reduced. The second row shows how the history length, $H$ influences $1-d_{s}$, where we can see that a larger value allows for a “smoother" transition. With the exception of small or large values, this parameter has the least effect on training. The third row illustrates how the leniency, $\lambda$ in Equation 5 and Algorithm 1 affects the training. Small values of $\lambda$ reduce noisy sources’ impacts influence earlier, whilst larger values miss some noisy sources all together. Here, the sources are coloured by their noise level, demonstrating that LAP reduced a source’s influence on training in a number of steps corresponding to a sources noise level.
@@
In our experiments, LAP parameters were set at $H=25$, $\delta=1.0$, $\lambda=0.8$ unless otherwise stated in Appendix A.7. These were chosen based on the synthetic experiments presented in Figure 2 and the validation results on CIFAR-10 (Appendix A.7).
4 Experiments
Baseline methods
To contextualise our approach, we evaluate varied baseline methods: (1) ARFL, designed to tackle label and input noise in a federated learning setting [17]; (2) a probabilistic method for instance dependent label noise (referred to here as IDPA) [15] which modifies the training loss; (3) “Co-teaching", which trains two models simultaneously to perform sample selection based on loss values during training [14]; (4) RRL, which uses contrastive learning and a $k$ nearest neighbours search to enforce a smoothness constraint on learnt representations [16], modifying model architecture and training loss; and (5) an identical model but without any specific modifications for tackling noise. Baseline methods were selected based on the availability of code, their use as baselines in the literature, and for variety in the methods used to evaluate the performance of LAP. Although our setting assumes noise is independent of features, we felt it was still beneficial to include IDPA, which is designed for instance dependent label noise. Additionally, for fair comparisons with federated learning, when models are compared to ARFL, they are trained on data that is structured such that each mini-batch contains a single source (in Appendix A.8 we study this restriction). This only concerns Table 1 since each batch in all other experiments contain multiple sources.
All baselines were implemented using the available code and trained using the recommended parameters with the model and data we test. For further details, see Appendix A.7.
Experimental design
To evaluate LAP, we employ various techniques to produce noisy data, extending the methods in Li et al. [17], Wang et al. [15], Han et al. [14], Li et al. [16] and test on nine datasets from computer vision, healthcare time-series, natural language processing, and tabular regression for diverse experiments.
@@ 4 Experiments
Table 1: Comparison of LAP with the baselines. Mean ± standard deviation of accuracy (%) of $10$ repeats of the baselines and LAP on synthetic data with different noisy types. For CIFAR-100 these numbers represent the top 5 accuracy. For CIFAR-10, CIFAR-100, and F-MNIST, the number of noisy sources are $4$, $2$, and $6$ out of $10$ respectively. Unreliable sources are each 100% noisy. Note there are non-significant differences in results on “Original Data" between Standard and LAP training since here the data is non-noisy.
Model Types
Noise Type Standard ARFL IDPA Co-teaching LAP (Ours)
CIFAR-10 Original Data 67.89 ± 1.16 61.32 ± 1.13 64.17 ± 1.92 67.19 ± 0.86 67.76 ± 1.61
Chunk Shuffle 63.78 ± 1.17 58.01 ± 1.54 58.11 ± 1.30 62.39 ± 0.65 64.22 ± 2.06
Random Label 57.47 ± 1.52 57.71 ± 2.32 51.02 ± 2.10 54.35 ± 1.48 62.74 ± 1.84
Batch Label Shuffle 57.48 ± 1.48 59.38 ± 1.25 51.71 ± 1.59 58.66 ± 0.67 63.06 ± 1.90
Batch Label Flip 51.83 ± 2.41 59.21 ± 1.35 49.58 ± 1.67 51.82 ± 1.41 63.21 ± 2.35
Added Noise 57.54 ± 1.48 51.57 ± 2.21 53.43 ± 2.14 57.64 ± 0.71 59.62 ± 1.91
Replace With Noise 61.11 ± 2.60 55.19 ± 1.72 59.22 ± 1.66 62.27 ± 1.02 61.52 ± 1.87
CIFAR-100 Original Data 64.46 ± 0.93 58.25 ± 1.19 58.64 ± 0.96 61.42 ± 0.99 64.46 ± 0.93
Chunk Shuffle 61.87 ± 1.16 57.12 ± 0.90 53.49 ± 1.22 54.74 ± 2.15 62.56 ± 0.61
Random Label 55.90 ± 1.11 56.49 ± 1.83 44.03 ± 1.31 45.08 ± 1.61 61.26 ± 1.02
Batch Label Shuffle 57.67 ± 0.86 56.85 ± 1.13 46.37 ± 1.17 50.32 ± 1.00 61.62 ± 1.34
Batch Label Flip 48.94 ± 2.13 56.34 ± 1.20 26.55 ± 7.71 24.97 ± 3.29 59.98 ± 1.06
Added Noise 56.83 ± 0.91 53.49 ± 1.01 48.35 ± 1.48 50.58 ± 0.99 59.56 ± 0.99
Replace With Noise 60.06 ± 0.97 55.00 ± 1.08 50.44 ± 2.24 53.00 ± 1.31 59.99 ± 1.26
F-MNIST Original Data 82.86 ± 1.51 80.95 ± 1.40 83.07 ± 1.06 77.85 ± 1.94 82.86 ± 1.51
Chunk Shuffle 76.90 ± 2.00 76.10 ± 1.15 76.82 ± 1.88 73.79 ± 3.12 81.03 ± 2.27
Random Label 73.47 ± 5.78 77.58 ± 1.69 74.30 ± 5.50 67.84 ± 6.00 77.91 ± 3.39
Batch Label Shuffle 81.26 ± 1.72 78.35 ± 1.17 81.43 ± 0.54 78.79 ± 0.55 81.53 ± 1.70
Batch Label Flip 77.64 ± 3.43 77.84 ± 1.50 78.22 ± 3.04 75.96 ± 2.83 79.26 ± 3.78
Added Noise 78.05 ± 1.96 73.12 ± 1.96 78.10 ± 1.69 75.26 ± 1.89 79.98 ± 3.83
Replace With Noise 78.99 ± 2.20 74.29 ± 1.19 79.09 ± 1.72 75.82 ± 2.14 81.14 ± 2.25
Following Li et al. [17], Wang et al. [15], Han et al. [14], Li et al. [16], we test our proposed method and baselines on CIFAR-10, CIFAR-100 [30], and F-MNIST [31] forming many comparisons with the literature. Along with Tiny-Imagenet and Imagenet [32], these five datasets form well-studied computer vision tasks, easing reproducibility. Additionally, we use a human labelled version of CIFAR-10, titled CIFAR-10N [33], for which we use the “worst labels" allowing us flexibility in our experiments. We then study an electrocardiograph (ECG) dataset, PTB-XL [34]; a time-series classification task with the goal to classify normal and abnormal cardiac rhythms and for which noise can be simulated following Wong et al. [35] and random labelling to understand LAP applied to time-series data with multiple noise types. A sentiment analysis natural language task allows us to further compare LAP against the literature, for which we employ the IMDB dataset [36] containing movie reviews and their sentiment. Finally, to illustrate our method’s use in regression tasks, we present results on the California Housing dataset [37]. For further information, see Appendix A.6.
To simulate data sources for CIFAR-10, CIFAR-100, F-MNIST, Imagenet, IMDB, and California Housing, data is uniformly split into $10$ distinct groups, and for Tiny-Imagenet we use $100$ groups to study larger numbers of sources. For CIFAR-10 and CIFAR-100, $4$ and $2$ of these sources are chosen to be noisy respectively; for F-MNIST, Imagenet, and Tiny-Imagenet $6$, $5$ and $40$ are chosen; and for IMDB and California Housing $4$ are chosen. These are in line with the noise rates used in the literature (often $20\%$, $40\%$, $50\%$). For CIFAR-10, in Appendix A.11 we also significantly increase the number of sources. To generate noise for the vision datasets, we extended the methods in Li et al. [17]: (1) Original Data: No noise is applied to the data; (2) Chunk Shuffle: Split features into distinct chunks and shuffle. This is only done on the first and/or second axis of a given input; (3) Random Label: Randomly assign a new label from the same code; (4) Batch Label Shuffle: For a given batch of features, randomly shuffle the labels; (5) Batch Label Flip: For a given batch of data, assign all features in this batch a label randomly chosen from the same batch; (6) Added Noise: Add Gaussian noise to the features, with mean $=0$ and standard deviation $=1$; (7) Replace with Noise: Replace features with Gaussian noise, with mean $=0$ and standard deviation $=1$.
For IMDB and California Housing we use random labelling, and for IMDB we further shuffle the order of text tokens (referred to as “Random Permute").
Although CIFAR-10N contains real-world noise, we must still split the data into sources. For each experiment, we assign sources as to evaluate varied levels of noise and numbers of noisy sources. As is done for PTB-XL, we linearly increase the number of noisy sources from $1$ to $7$ (out of $10$ in total), and for each set of noisy sources we linearly increase the noise level from $0.25$ to $1.0$.
For this, five base model architectures are evaluated (of multiple sizes): A Multilayer Perceptron, Convolutional Neural Network, a 1D and 2D ResNet [38], and an LSTM [39] (Appendix A.7).
In all experiments, data points contain an observation, source, and target, which are assigned to mini-batches in the ordinary way. Features and labels are passed to the model for training, whilst sources are used by LAP (Appendix A.7). The test sets contain clean labels only.
Results
@@
Noise Type RRL RRL + LAP (Ours)
@@
Baselines and LAP. Table 1 shows the results of the baselines and LAP on CIFAR-10, CIFAR-100, and F-MNIST.
These experiments illustrate that in high noise settings, in most cases, using LAP allows for improved test performance over the baselines. In particular, our method allows for improved performance over two different approaches to learn from noisy labels (IDPA and Co-teaching), and over an approach from federated learning (ARFL), as well as training a model in the standard way. This is the case across the three different datasets tested in this experiment with varied noise types, displaying LAP’s wide applicability. Required by ARFL, each batch contains one source and so within Appendix A.8 we replicate the CIFAR-10 experiments with multiple sources in each batch. There we see that all methods improve on the results shown in Table 1 by approximately the same amount, leading to the same conclusions.
LAP in conjunction with RRL. When evaluating the RRL baseline, we found that using the same model architecture as in Table 1 caused an unexpected loss in accuracy. We therefore used the larger architecture (Appendix A.7) employed in Li et al. [16] and data augmentation to allow for a fairer comparison. We tested RRL with and without LAP to evaluate its ability to be used in conjunction with other methods. Table 2 presents these results on CIFAR-10, and further underlines the use of LAP to improve model performance on data generated by multiple sources. Here, LAP supplements the performance obtained by RRL, which uses a contrastive learning approach to tackle noisy features and labels, that is improved upon by further applying our method to make use of the source values. We observe that LAP increases all metric scores except for “Replace With Noise", where the difference is relatively small, and “Original Data", where both methods perform equally. In Appendix A.9, we further experiment with different numbers of noisy sources and noise rates in the same setting.
@@
Figure 3: LAP performance with a varied number of sources. In 3(a) we show the area under the precision-recall curve for standard training and using LAP on PTB-XL with label noise and simulated ECG interference noise for $12$ total sources. In 3(b) we show the accuracy on CIFAR-10N with real human labelling noise when using RRL and RRL + LAP, with $10$ sources. In both, the noise of the sources increases linearly from $25\%$ to $100\%$ for each number of noisy sources. The lines and error bands represent the mean and standard deviation of the maximum value for each of the 5 repeats. These figures illustrate that LAP maintains higher performance as noise rates increase.
Table 3: Standard and LAP results on a natural language task. Mean ± standard deviation of maximum test accuracy (%) of $5$ repeats of standard training and LAP on the IMDB dataset with different types of noise. Here, $4$ out of $10$ sources are 100% noisy. Note that “Random Label" is the noise type with a largest difference between methods.

### v4

@@ 4 Methods
Figure 3: Visualisation of Equation 4. Each colour represents the loss values from a single source over a small number of steps, with its density weighted by its temperature, $w(s)=f(C_{s})$. This shows how sources contribute to $\mu_{s^{\prime}}$ and $\sigma_{s^{\prime}}^{2}$ as their $C_{s}$ changes during training and given the leniency $\lambda$. These values are synthetic and for demonstration.
Note that in this construction, we make no assumptions about the type of observations contained in $\mathcal{D}$. Therefore, our method is applicable to images, natural language, time series, and tabular data for both regression and classification tasks.
5 Results
Baseline methods
To contextualise our approach, we evaluate varied baseline methods: (1) ARFL, designed to tackle label and input noise in a federated learning setting [12]; (2) IDPA, a probabilistic method for instance dependent label noise [16] which modifies the training loss; (3) “Co-teaching”, which trains two models simultaneously to perform sample selection based on loss values during training [15]; (4) CDR, which regularises model weights such that important parameters, learnt during the early training phase, do not considerably change when the model fits to the noisy data [14]; (5) Label Smoothing, which uses a combination of the true label and a uniformly weighted soft label during training [18]; (6) RRL, which uses contrastive learning and a $k$ nearest neighbours search to enforce a smoothness constraint on learnt representations [17], modifying model architecture and training loss; and (7) an identical model but without any specific modifications for tackling noise. Baseline methods were selected based on the availability of code, their use as baselines in the literature, and for variety in the methods used to evaluate the performance of LAP. Although our setting assumes noise is independent of features, we felt it was still beneficial to include IDPA, which is designed for instance dependent label noise. Note that because ARFL is a federated learning approach, data cannot be shuffled in the same way as the other models, since each client trains on a single source.
It is important to note that both IDPA and Co-teaching require the training of a model twice, making LAP significantly less computationally expensive.
All baselines are implemented using the available code and trained using the recommended parameters with the model and data we test. For further details, see Appendix A.9.
Experimental design
To evaluate LAP, we employ various techniques to produce noisy data, extending the methods in Li et al. [12], Wang et al. [16], Han et al. [15], Li et al. [17] and test on eleven datasets from computer vision, healthcare time-series, natural language processing, and tabular regression for diverse experiments.
Following Li et al. [12], Wang et al. [16], Han et al. [15], Li et al. [17], we test our proposed method and baselines on CIFAR-10, CIFAR-100 [32], and F-MNIST [33] forming many comparisons with the literature. Along with Tiny-Imagenet and Imagenet [34], these five datasets form well-studied computer vision tasks, easing reproducibility. Additionally, we use a human-labelled version of CIFAR-10, titled CIFAR-10N [35], for which we use the “worst labels” allowing us flexibility in our experiments. We then study an electrocardiograph (ECG) dataset, PTB-XL [36]; a time-series classification task with the goal of classifying normal and abnormal cardiac rhythms and for which noise can be simulated following Wong et al. [37] and random labelling to understand LAP applied to time-series data with multiple types of noise. Additionally, a natural language sentiment analysis task allows us to compare LAP with the literature, for which we use the IMDB data set [38] that contains movie reviews and their sentiment. Next, we use the GoEmotions dataset [39], a natural language emotion prediction task, which contains unbalanced source sizes and class distributions in the real world, allowing us to test the robustness of LAP to varied source constructions. Finally, to illustrate the use of our method for regression, we present results on the California Housing dataset [40]. For further information, see Appendices A.8 and A.15.
To simulate data sources for CIFAR-10, CIFAR-100, F-MNIST, Imagenet, IMDB, and California Housing, data is uniformly split into $10$ distinct groups, and for Tiny-Imagenet we use $100$ groups to study larger numbers of sources. For CIFAR-10 and CIFAR-100, $4$ and $2$ of these sources are chosen to be noisy respectively; for F-MNIST, Imagenet, and Tiny-Imagenet $6$, $5$ and $40$ are chosen; and for IMDB and California Housing $4$ are chosen. These are in line with the noise rates used in the literature (often $20\%$, $40\%$, $50\%$). For CIFAR-10, in Appendix A.14 we also significantly increase the number of sources. To generate noise for the vision datasets, we extended the methods in Li et al. [12]: (1) Original Data: No noise is applied to the data; (2) Chunk Shuffle: Split features into distinct chunks and shuffle. This is only done on the first and/or second axes of a given input; (3) Random Label: Randomly assign a new label from the same code; (4) Batch Label Shuffle: For a given batch of features, randomly shuffle the labels; (5) Batch Label Flip: For a given batch of data, assign all features in this batch a label randomly chosen from the same batch; (6) Added Noise: Add Gaussian noise to the features, with mean $=0$ and standard deviation $=1$; (7) Replace with Noise: Replace features with Gaussian noise, with mean $=0$ and standard deviation $=1$.
For IMDB, GoEmotions, and California Housing, we use random labelling.
Since GoEmotions is annotated by $82$ raters, their identification number is used to form the data sources. Some raters contributed a handful of data points, whilst others labelled thousands; with each rater providing a different distribution of class labels. This allowed us to explore the real-world case of data with sources of imbalanced size and label distribution. A detailed discussion of the imbalanced sources is given in Appendix A.15.
Although CIFAR-10N contains real-world noise, we still have to split the data into sources. For each experiment, we assign sources to evaluate the varied levels of noise and the number of noisy sources. As in PTB-XL, we linearly increase the number of noisy sources from $1$ to $7$ (out of $10$ in total), and for each set of noisy sources, we linearly increase the noise level from $0.25$ to $1.0$.
For this, seven base model architectures (of multiple sizes) are evaluated: A Multilayer Perceptron, Convolutional Neural Networks, a 1D and 2D ResNet [41], an LSTM [42], and a transformer encoder [43] (Appendix A.9).
In all experiments, data points contain an observation, source, and target, which are assigned to mini-batches in the ordinary way. Features and labels are passed to the model for training, whilst sources are used by LAP (Appendix A.9). The test sets contain only clean observations.
Intuition for the hyperparameters
@@ 5 Results
Figure 3 shows how the training process evolves over time when using LAP. At first, all sources are considered equally reliable (that is, $T_{s}=1$), and therefore the weighted mean of the loss values is the mean of all sources. As the source temperature changes, the weighed mean plus $\lambda$ standard deviations of the losses moves toward the sources with lower expected losses, allowing more learning from these over time.
@@
Figure 4 illustrates how $1-d_{s}$ (the temperature) in Equation 5 develops during training for each source and different values of $H$, $\delta$, and $\lambda$. The first row shows how varying the depression strength, $\delta$ affects the rate at which the learning on a source is reduced. The second row shows how the history length, $H$ influences $1-d_{s}$, where we can see that a larger value allows for a “smoother” transition. With the exception of small or large values, this parameter has the least effect on training. The third row illustrates how the leniency $\lambda$ in Equation 4 and Algorithm 1 affects training. Small values of $\lambda$ reduce the influence of noisy sources earlier, whilst larger values miss some noisy sources altogether. Here, the sources are coloured by their noise level, demonstrating that LAP reduced the influence of a source on training in a number of steps corresponding to the noise level of the source.
@@
In our experiments, the parameters of LAP were set at $H=25$, $\delta=1.0$, $\lambda=0.8$ unless otherwise stated in Appendix A.9. These were chosen based on the synthetic experiments presented in Figure 4 and the discussion of the effects of these hyperparameters provided in Appendix A.10.
LAP and the baselines
@@
Table 1: Comparison of LAP with the baselines. Mean ± standard deviation of the maximum test accuracy (%) of $5$ repeats of the baselines and LAP on synthetic data with different noisy types. For CIFAR-100 these numbers represent the top 5 accuracy. For CIFAR-10, CIFAR-100, and F-MNIST, the number of noisy sources are $4$, $2$, and $6$ out of $10$ respectively. Unreliable sources are each 100% noisy. All values in bold are within 1 standard deviation of the maximum score. Note that IDPA and Co-teaching require almost twice the training time in comparison to Standard, ARFL, CDR, Label Smoothing (LS), and LAP.
Model Types Noise Type Standard ARFL (2022) IDPA (2021) Co-teaching (2018) CDR (2021) LS (2022a) LAP (Ours) CIFAR-10 Original Data 77.82 ± 0.37 74.89 ± 1.67 79.89 ± 1.01 80.04 ± 0.49 77.2 ± 1.59 78.47 ± 0.72 77.52 ± 0.9 Chunk Shuffle 72.12 ± 0.95 68.42 ± 0.71 72.68 ± 0.96 74.77 ± 0.3 70.02 ± 1.35 70.1 ± 0.97 74.29 ± 1.43 Random Label 67.84 ± 1.84 69.93 ± 1.88 55.66 ± 1.54 68.61 ± 1.78 65.61 ± 0.97 67.92 ± 2.66 73.44 ± 0.9 Batch Label Shuffle 68.75 ± 1.36 67.12 ± 1.93 68.19 ± 1.48 71.68 ± 0.92 68.21 ± 1.06 67.35 ± 1.23 73.94 ± 0.33 Batch Label Flip 67.12 ± 0.7 66.16 ± 1.96 69.65 ± 1.54 71.18 ± 0.76 66.59 ± 1.79 68.39 ± 2.14 73.85 ± 0.92 Added Noise 70.23 ± 0.76 68.59 ± 1.38 70.94 ± 1.61 72.29 ± 0.61 68.3 ± 1.32 69.48 ± 1.33 72.63 ± 1.24 Replace With Noise 72.79 ± 0.67 67.01 ± 1.02 71.96 ± 1.53 73.87 ± 0.57 71.02 ± 1.04 72.97 ± 0.83 73.84 ± 0.55 CIFAR-100 Original Data 75.95 ± 1.01 60.42 ± 1.91 77.78 ± 0.95 76.25 ± 0.54 75.91 ± 1.11 75.42 ± 0.49 76.01 ± 0.54 Chunk Shuffle 68.86 ± 0.74 56.76 ± 2.69 66.82 ± 0.85 69.58 ± 0.51 68.65 ± 0.91 68.15 ± 0.95 70.41 ± 0.96 Random Label 58.34 ± 0.91 48.85 ± 3.2 49.5 ± 1.03 61.2 ± 0.54 57.12 ± 0.76 58.3 ± 0.93 69.05 ± 0.38 Batch Label Shuffle 64.8 ± 0.97 58.2 ± 4.45 64.34 ± 2.09 69.34 ± 0.69 65.69 ± 1.15 64.73 ± 1.12 69.31 ± 1.4 Batch Label Flip 61.89 ± 0.97 56.51 ± 3.36 64.61 ± 2.1 69.04 ± 1.19 60.01 ± 1.16 60.64 ± 2.05 69.82 ± 0.78 Added Noise 65.28 ± 0.77 58.35 ± 4.16 64.44 ± 1.11 66.02 ± 0.36 65.24 ± 1.32 65.38 ± 1.28 67.69 ± 1.17 Replace With Noise 67.76 ± 1.33 58.88 ± 4.33 66.06 ± 0.42 68.52 ± 1.31 68.48 ± 1.31 67.58 ± 0.92 68.71 ± 1.22 F-MNIST Original Data 83.64 ± 0.33 82.0 ± 0.4 83.69 ± 0.65 79.09 ± 1.14 82.42 ± 2.27 83.74 ± 0.27 83.52 ± 0.21 Chunk Shuffle 77.12 ± 1.79 77.69 ± 0.51 77.45 ± 2.21 74.74 ± 1.21 76.45 ± 1.7 78.17 ± 0.75 82.18 ± 1.26 Random Label 76.22 ± 8.06 77.31 ± 3.82 76.57 ± 7.16 77.41 ± 4.47 75.33 ± 7.27 74.79 ± 4.84 81.8 ± 1.37 Batch Label Shuffle 82.65 ± 0.13 79.11 ± 1.71 82.82 ± 0.43 82.3 ± 0.32 82.88 ± 0.44 82.63 ± 0.45 82.45 ± 0.38 Batch Label Flip 80.25 ± 1.87 78.85 ± 1.78 82.31 ± 0.27 80.43 ± 0.81 81.29 ± 0.58 80.56 ± 0.67 79.87 ± 0.7 Added Noise 77.88 ± 1.86 73.28 ± 1.31 78.42 ± 2.1 75.72 ± 1.55 76.21 ± 2.37 79.02 ± 0.75 78.72 ± 2.02 Replace With Noise 79.77 ± 0.8 74.12 ± 1.36 80.64 ± 0.42 78.76 ± 1.06 78.24 ± 0.93 79.94 ± 0.64 82.94 ± 0.31
Table 1 shows the results of the baselines and LAP on CIFAR-10, CIFAR-100 and F-MNIST with varied noise. This illustrates that in a high noise setting, LAP often allows for improved test accuracy over the baselines, four different approaches to learn from noisy labels (IDPA, Co-teaching, CDR, and Label Smoothing), an approach from federated learning (ARFL), and training a model in the standard way. Interestingly, IDPA and Co-teaching achieve a higher accuracy on the original data in CIFAR-10 and 100, likely because IDPA and Co-teaching require the training of a model for twice as long or twice, almost doubling training time over LAP, ARFL and standard training.
On CIFAR-10 and 100, the performance improvement with LAP is most apparent with the random labelling noise. For example, on CIFAR-100, LAP achieves a $+18.4\%$ increase in top-5 accuracy over standard training, whilst IDPA performs worse. Although the performance improvements of LAP are less dramatic on F-MNIST due to the simpler nature of the dataset, the model still demonstrates better accuracy in certain noise conditions, particularly with chunk shuffling and random label noise.
In these experiments, LAP achieves the best accuracy or is within a few percentage points of the best accuracy. In the latter case, the best accuracy is then often achieved by a method that requires longer training but that is less robust to other noise types. Therefore, LAP overall achieves consistently higher accuracy while requiring less computation.
To further illustrate the potential accuracy improvement from using LAP, in Appendix A.11 we also present these results as a percentage difference from the standard training method. In the following more challenging experiments, the improvement in the accuracy of using LAP becomes clearer. Furthermore, although IDPA and co-teaching are limited to classification tasks, LAP also works on regression (Appendix A.19).
LAP in conjunction with RRL
Since RRL requires significant changes to the model architecture and data augmentation, we separately test the large architecture (Appendix A.9) and the experimental setup used in Li et al. [17] for a better comparison. Here, we test RRL with and without LAP to evaluate its ability to be used in conjunction with other methods. Table 2 presents these results on CIFAR-10, and further highlights the use of LAP to improve model performance on data generated from multiple sources. Here, LAP supplements the accuracy obtained by RRL, which uses a contrastive learning approach to tackle noisy features and labels that is improved upon by further applying our method to make use of the source values. We observe that LAP increases all metric scores except for “Replace With Noise”, where the difference is relatively small, and “Original Data”, where both methods perform equally. Again, the most substantial increase in accuracy comes from random labelling noise, where LAP improves RRL by approximately $+5\%$. These results are supplemented by Appendix A.12, where we test different numbers of noisy sources and noise rates.
@@
Noise Type RRL (2021) RRL + LAP (Ours)
@@
Varied noise level and source sizes with real-world time-series data
Figure 5(a) compares the results of not using and using LAP for a time-series classification task on normal and abnormal cardiac rhythms with both label and input noise. To evaluate the model on data sources with differing noise levels, for each number of noisy sources we set the noise levels of the sources in linear increments between $25\%$ and $100\%$. Additionally, we linearly increase the number of these noisy sources and measure the area under the precision-recall curve (AUC PR) on the test set. These results show that LAP allows for improved model training on data with sources of varied noise and with multiple noise types. Figure 5(a) also illustrates that in this setting, LAP is robust to increases in noise within the dataset, as AUC PR does not degrade significantly, especially when compared to not using LAP. Furthermore, the standard deviation of the results is smaller, suggesting that LAP is more consistent in its AUC-PR. When only two sources contain noise, the total noise rate in the dataset is only $10\%$, which has little negative effect on standard training.
@@
Figure 5: LAP results with a varied number of sources and noise levels. In 5(a) we show the area under the precision-recall curve for standard training and using LAP on PTB-XL with label noise and simulated ECG interference noise for $12$ total sources. In 5(b) we show the accuracy on CIFAR-10N with real human labelling noise when using RRL and RRL + LAP, with $10$ sources. In both, the noise of the sources varies linearly from $25\%$ to $100\%$ for each number of noisy sources. The lines and error bands represent the mean and standard deviation of the maximum value for each of the 5 repeats. These figures illustrate that LAP maintains higher performance as noise rates increase.
Real-world noisy labels
To further demonstrate the effectiveness of LAP in conjunction with other noisy data methods and on real-world noise, we now present the results on a dataset with human labelling of varied noise rates. Figure 5(b) illustrates the results on CIFAR-10N, a noisy human-labelled version of CIFAR-10, where we observe a pattern similar to that depicted in Figure 5(a). In both cases, LAP improved the accuracy on noisy data at all noise levels tested, and the standard deviation of the accuracy is reduced. However, we also note that in this case the accuracy of LAP degraded faster than the AUC PR in Figure 5(a) as the number of noisy sources increased. This could be for various reasons, such as differing noise types, model baselines, and data type. In Appendix A.13 we also show the results of this experiment when using the same model architecture as in Table 1, and with standard training as a baseline, instead of RRL, where we arrive at the same conclusions.
LAP on a noisy natural language task
Table 3: Baselines and LAP results on a natural language task. Mean ± standard deviation of maximum test accuracy (%) of $5$ repeats of the methods on the IMDB dataset with different types of noise. Here, $4$ out of $10$ sources are 100% noisy.
Model Types Noise Type Standard IDPA (2021) Co-teaching (2018) CDR (2021) LS (2022a) LAP (Ours) Original Data 82.81 ± 0.79 83.24 ± 0.56 85.12 ± 0.8 82.49 ± 1.49 83.89 ± 0.91 83.2 ± 0.83 Random Label 65.01 ± 0.65 64.61 ± 0.71 67.18 ± 0.69 64.8 ± 0.91 66.34 ± 1.13 71.95 ± 2.94
We also tested LAP on a natural language task, in which the goal is to predict the sentiment from the review of the movie. The results are presented in Table 3, for which we tested two types of noise. Here, LAP clearly outperforms the baselines by a significant margin, with IDPA and CDR not improving standard training. Additionally, we expect that when the data is not noisy, using LAP should not limit performance, as we see here. However, it is interesting that Co-teaching performs better in the “Original Data”. We hypothesise that the IMDB dataset contains some noisy labels (through human error), which were split uniformly across data sources; a limitation of our method that we discuss in Section 6.
Imagenet results with multiple noise types
We additionally present results on Imagenet, which allow us to test LAP on a large-scale dataset. As with the experiment in Figure 5(a), noisy sources on Imagenet contain both input and label noise. In Figure 5(a) all noisy sources contain two types of noise, however in the experiment presented in Table 4, we allow different sources to contain different types of noise. These results further illustrate that the use of LAP improves model performance when training data is noisy, since it achieves a greater maximum top-5 test accuracy than the standard training method. In Appendix A.18, we also study the learning curve on this dataset to understand the effects of early-stopping used in combination with our method.
@@
Noise Type Standard LAP (Ours)
Input and Label Noise 68.05 ± 0.26 70.61 ± 0.26
Real-world and imbalanced data sources
Table 5: Results on GoEmotions, an imbalanced sources dataset. Mean ± standard deviation of maximum test top-5 accuracy (%) over $5$ repeats of the methods on the GoEmotions data with noisy labels and random sentence permutation. Here, $30$ out of $82$ sources are 100% noisy.